# Knowledge-Centric Templatic Views of Documents

### v1

@@ A.1 Questions
The above criteria are non-exhaustive. Not all criteria must be met, and you may use other relevant criteria to make your decision. You are not rating the document for factual correctness, and only need to refer to the corresponding scientific article if it will aid in making your preference. You can answer this question with either Document A or Document B.
@@ A.2 Edge cases
Template Temp R-1 R-L BertScore
slides 0 7.7 7.3 35.1
0.25 7.8 7.2 35.4
0.5 7.6 7.0 36.4
0.75 7.5 7.0 35.5
1 8.0 7.4 35.8
@@
0.25 14.2 13.0 53.1
0.5 13.6 12.5 52.2
0.75 13.2 12.0 50.8
1 12.6 11.6 50.3
Table 5: Results of the temperature hyperparameter experiments. There seems to be little consistency in performance as related to temperature across different templates.
Appendix B Temperature Experiments
We experiment with the temperature of the generations to see how temperature affects performance. Similar to the textual representation experiments, described in §6.1.1, we randomly sample 100 documents each from LongSumm and Doc2PPT for the blog and poster generation tasks, respectively. We use the entirety of the Paper-Poster dataset, since it contains less than 100 examples. We use gpt35-16k and experiment with the temperatures $[0.0,0.25,0.5,0.75,1.0]$. The results of this experiment can be found in Table 5. As we can see from the results, there seems to be little consistency across the different types in which temperature performs the best.

### v2

@@ B.1 Questions
The above criteria are non-exhaustive. Not all criteria must be met, and you may use other relevant criteria to make your decision. You are not rating the document for factual correctness, [^6] The annotators are non-experts and do not have the background to determine factual correctness of scientific information. Instead, they are encouraged to use the original paper to understand if the information presented in the documents represent the information in the paper, to the best of their understanding. and only need to refer to the corresponding scientific article if it will aid in making your preference. You can answer this question with either Document A or Document B.
@@ B.2 Edge cases
Simularity Measure
Temp R-L M B BERTS
Slides 0.0 7.3 8.2 1.6 35.1
0.25 7.2 8.3 1.5 35.4
0.5 7.0 8.3 1.5 36.4
0.75 7.0 8.0 1.2 35.5
1.0 7.4 8.2 1.4 35.8
@@
0.25 13.0 14.8 3.4 53.1
0.5 12.5 14.0 2.7 52.2
0.75 12.0 13.9 3.0 50.8
1.0 11.6 11.9 2.4 50.3
Table 5: Results of the temperature hyperparameter experiments. We use ROUGE-L (R-L), METEOR (M), BLEU (B) and BERTScore (BERTS) as our similarity measures.
Appendix C Temperature Experiments
We experiment with the temperature of the generations to see how temperature affects performance. We randomly sample 100 documents each from LongSumm and Doc2PPT for the blog and poster generation tasks, respectively. We use the entirety of the Paper-Poster dataset, since it contains less than 100 examples. We use gpt35-16k and experiment with the temperatures $[0.0,0.25,0.5,0.75,1.0]$. The results of this experiment can be found in Table 5. As we can see from the results, there seems to be little consistency across the different types in which temperature performs the best.
Appendix D Longsumm Blind Test Set Results
We submit the final documents from GPT4, the best performing model overall, to the Longsumm blind test set evaluation. We compare the documents generated with and without the intermediate step. We see that without the intermediate representation we get a Rouge-1 score of 46.8 while the results generated without the intermediate representation received a Rouge-1 score of 46.4. We note that this blind test set of 22 papers is significantly smaller than the evaluation data (505 papers) we used in the main body of this paper. Despite not designing a task specific method, we place second on the leaderboard, showing the powerful capabilities of LLMs in long document generation.
Appendix E Example Outputs
We provide examples of the outputs generated with and without the intermediate representation below. The documents in all examples are generated with GPT4 (gpt4-32k). Figure 5 includes example slide generations, Figure 6 includes example blog generations, and Figure 7 includes example poster generations.
(a) Document generated without intermediate representation. This example is not cropped.
(b) Document generated with intermediate representation. This example is cropped for space and includes an additional 4 slides that are not included for space.
Figure 5: The above documents are example slides generated by GPT4 (gpt4-32k) with and without the intermediate representation. We can see that without the intermediate step, the model did not generate a true slide deck.
(a) Document generated without the intermediate representation. This example is not cropped.
(b) Document generated with the intermediate representation. This example is cropped for space and includes an additional page of text.
Figure 6: The above documents are example blog posts generated by GPT4 (gpt4-32k) with and without the intermediate representation. We can see that without the intermediate representation, the model did not properly format the LaTeX file for compilation.
(a) Document generated without the intermediate representation. This example is cropped for space and includes an additional 3 slides.
(b) Document generated with the intermediate representation. This example is cropped for space and includes an additional 4 slides.
Figure 7: The above documents are example posters generated by GPT4 (gpt4-32k) with and without the intermediate representation. We found that GPT4 often generates slide decks in place of posters. We can see that the document generated without the intermediate representation contains more verbose panels and includes less formatting.